(1 point) Find the solution of x²y + 5xy + (4 – 3x) y = 0, x > 0 of the form y=x Wazek, k=0 where ao = 1. r = help (number
Answer #1

Given. + 5 ry +_14-3x Sy= 0. Hisari 00 ciletei). 06 x K44-1 - 2 ak SCK+1) CIT87) 1920 Substituting ini ! differential @gon.Tags الالاDate Ifor neo 1()» (x+4) +47 Coreet E kal Tokrr) (k+4+4 9+4 ] ac u 175 ok+x=0 .3 Eakat they Recto indicial ! egin.. Page Date And y, 2 = {(1121 x 21h-2) .. ko lng. Y 2 = { x 21h-2) no (n!)